# Deflated Dynamics Value Iteration

### v1

@@ Abstract
Abstract: The Value Iteration (VI) algorithm is an iterative procedure to compute the value function of a Markov decision process, and is the basis of many reinforcement learning (RL) algorithms as well. As the error convergence rate of VI as a function of iteration $k$ is $O(\gamma^{k})$, it is slow when the discount factor $\gamma$ is close to $1$. To accelerate the computation of the value function, we propose Deflated Dynamics Value Iteration (DDVI). DDVI uses matrix splitting and matrix deflation techniques to effectively remove (deflate) the top $s$ dominant eigen-structure of the transition matrix ${\mathcal{P}}^{\pi}$. We prove that this leads to a $\tilde{O}(\gamma^{k}|\lambda_{s+1}|^{k})$ convergence rate, where $\lambda_{s+1}$is $(s+1)$-th largest eigenvalue of the dynamics matrix. We then extend DDVI to the RL setting and present Deflated Dynamics Temporal Difference (DDTD) algorithm. We empirically show the effectiveness of the proposed algorithms.
Deflated Dynamics Value Iteration
Jongmin Lee1, Amin Rakhsha2, Ernest K. Ryu3, Amir-massoud Farahmand2
1 Seoul National University, South Korea
2 University of Toronto & Vector Institute, Canada
3 University of California, Los Angeles, USA
\@abstract
1.  Introduction
Computing the value function $V^{\pi}$ for a policy $\pi$ or the optimal value function $V^{\star}$ is an integral step of many planning and reinforcement learning (RL) algorithms. Value Iteration (VI) is a fundamental dynamic programming algorithm for computing the value functions, and its approximate and sample-based variants, such as Temporal Different Learning [Sutton, 1988], Fitted Value Iteration [Ernst et al., 2005, Munos and Szepesvári, 2008], Deep Q-Network [Mnih et al., 2015], are the workhorses of modern RL algorithms [Bertsekas and Tsitsiklis, 1996, Sutton and Barto, 2018, Szepesvári, 2010, Meyn, 2022].
@@ 1.  Introduction
Recently, there has been a growing body of research that explores the application of acceleration techniques of other areas of applied math to planning and RL: Geist and Scherrer [2018], Sun et al. [2021], Ermis and Yang [2020], Park et al. [2022], Ermis et al. [2021], Shi et al. [2019] applies Anderson acceleration of fixed-point iterations, Lee and Ryu [2023] applies Anchor acceleration of minimax optimization, Vieillard et al. [2020], Goyal and Grand-Clément [2022], Grand-Clément [2021], Bowen et al. [2021], Akian et al. [2022] applies Nesterov acceleration of convex optimization and Farahmand and Ghavamzadeh [2021] applies ideas inspired by PID controllers in control theory.
More detail is provided in Appendix 7.1.
We introduce a novel approach to accelerate VI based on modification of the eigenvalues of the transition dynamics, which are closely related to the convergence of VI. To see this connection, consider the policy evaluation problem where the goal is to find the value function $V^{\pi}$ for a given policy $\pi$. VI starts from an arbitrary $V^{0}$ and iteratively sets $V^{k+1}\leftarrow r^{\pi}+\gamma{\mathcal{P}}^{\pi}V^{k}$, where $r^{\pi}$ and ${\mathcal{P}}^{\pi}$ are the reward vector and the transition matrix of policy $\pi$, respectively. If $V^{0}=0$, the error vector at iteration $k$ can be shown to be $V^{\pi}-V^{k}=\sum_{i=k}^{\infty}(\gamma{\mathcal{P}}^{\pi})^{i}r^{\pi}$. Let us take a closer look at this difference.
@@
The diagonal terms are of the $(\gamma\lambda_{i})^{k}$ form, so they all converge to zero. The dominant term is $(\gamma\lambda_{1})^{k}$, which corresponds to the largest eigenvalue. As the largest eigenvalue $\lambda_{1}$ of the stochastic matrix ${\mathcal{P}}^{\pi}$ is $1$, this leads to the dominant behaviour of $O(\gamma^{k})$, which we also get from the norm-based contraction mapping analysis. The second dominant term behaves as $O((\gamma|\lambda_{2}|)^{k})$, and so on.
If we could somehow remove from ${\mathcal{P}}^{\pi}$ the subspace corresponding to the top $s$ eigen-structure with eigen-values $\lambda_{1},\dotsc,\lambda_{s}$, the dominant behaviour of the new procedure would be $O((\gamma|\lambda_{s+1}|)^{k})$, which can be much faster than $O(\gamma^{k})$ of the conventional VI. Although this is perhaps too good to seem feasible, this is exactly what the proposed Deflated Dynamics Value Iteration (DDVI) algorithm achieves. Even if $\mathcal{P}^{\pi}$ is not a diagonalizable matrix, DDVI works. DDVI is based on two main ideas.
The first idea is the deflation technique, well-studied in linear algebra (see Section 4.2 of Saad [2011]), that allows us to remove large eigenvalues of ${\mathcal{P}}^{\pi}$. This is done by subtracting a matrix $E$ from ${\mathcal{P}}^{\pi}$. This gives us a “deflated dynamics” ${\mathcal{P}}^{\pi}-E$ that does not have eigenvalues $\lambda_{1},\dotsc,\lambda_{s}$. The second idea is based on matrix splitting (see Section 11.2 of Golub and Van Loan [2013]), which allows us to use the modified dynamics ${\mathcal{P}}^{\pi}-E$ to define a VI-like iterative procedure and still converge to the same solution $V^{\pi}$ to which the conventional VI converges.
Deflation has been applied to various algorithms such as conjugate gradient algorithm [Saad et al., 2000], principal component analysis [Mackey, 2008], generalized minimal residual method [Morgan, 1995], and nonlinear fixed point iteration [Shroff and Keller, 1993] for improvement of convergence. While some prior in accelerated planning can be considered as special cases of deflation [Bertsekas, 1995, White, 1963], to the best of our knowledge, this is the first application of the deflation technique that eliminates multiple eigenvalues in the context of RL. On the other hand, multiple planning algorithms have been introduced based on the general idea of matrix splitting [Hastings, 1968, Kushner and Kleinman, 1968, 1971, Reetz, 1973, Porteus, 1975, Rakhsha et al., 2022], though with a different splitting than this work.
@@
2.  Background
We first briefly review basic definitions and concepts of Markov Decision Processes (MDP) and Reinforcement Learning (RL) [Bertsekas and Tsitsiklis, 1996, Sutton and Barto, 2018, Szepesvári, 2010, Meyn, 2022]. We then describe the Power Iteration and QR Iterations, which can be used to compute the eigenvalues of a matrix.
@@ 2.  Background
The discounted MDP is defined by the tuple $(\mathcal{X},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{X}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}\colon\mathcal{X}\times\mathcal{A}\rightarrow\mathcal{M}(\mathcal{X})$ is the transition probability kernel, $\mathcal{R}\colon\mathcal{X}\times\mathcal{A}\rightarrow\mathcal{M}()$ is the reward kernel, and $\gamma\in[0,1)$ is the discount factor. In this work, we assume that the MDP has a finite number of states. We use $r\colon\mathcal{X}\times\mathcal{A}\rightarrow$ to denote the expected reward at a given state-action pair. Denote $\pi\colon\mathcal{X}\rightarrow\mathcal{M}(\mathcal{A})$ for a policy. We define the reward function for a policy $\pi$ as $r^{\pi}(x)=\mathbf{E}_{a\sim\pi(\cdot\mid x)}\left[r(x,a)\right]$. The transition kernel of following policy $\pi$ is denoted by ${\mathcal{P}}^{\pi}\colon\mathcal{X}\rightarrow\mathcal{M}(\mathcal{X})$ and is defined as
${\mathcal{P}}^{\pi}(x\rightarrow x^{\prime})$ $=\mathrm{Prob}(x\rightarrow x^{\prime}\,|\,a\sim\pi(\cdot\mid x),x^{\prime}\sim\mathcal{P}(\cdot\mid x,a)).$
The value function for a policy $\pi$ is $V^{\pi}(x)=\mathbf{E}_{\pi}[\sum^{\infty}_{t=0}\gamma^{t}r(X_{k},a_{t})\mid x_{0}=x]$ where $\mathbf{E}_{\pi}$ denotes the expected value over all trajectories $(x_{0},a_{0},x_{1},a_{1},\dots)$ induced by $\mathcal{P}$ and $\pi$. We say that $V^{\star}$ is optimal value functions if $V^{\star}=\sup_{\pi}V^{\pi}$. We say $\pi^{\star}$ is optimal policies if $\pi^{\star}=\operatorname*{argmax}_{\pi}{V^{\pi}}$.
The Bellman operator $T^{\pi}$ for policy $\pi$ is defined as the mapping that takes $V\in\mathcal{F}(\mathcal{X})$ and returns a new function such that its value at state $x$ is $(T^{\pi}V)(x)=\mathbf{E}_{a\sim\pi(\cdot\mid x),x^{\prime}\sim\mathcal{P}(\cdot\mid x,a)}\left[r(x,a)+\gamma V(x^{\prime})\right]$. The Bellman optimality operator ${T^{\star}}$ is defined as ${T^{\star}}V(x)=\sup_{a\in\mathcal{A}}\left\{r(x,a)+\gamma\mathbf{E}_{x^{\prime}\sim\mathcal{P}(\cdot\,|\,x,a)}\left[V(x^{\prime})\right]\right\}$. The value functions $V^{\pi}$ and $V^{\star}$ are the fixed points of the Bellman operators, that is, $V^{\pi}=T^{\pi}V^{\pi}$ and $V^{\star}=T^{\pi}V^{\star}$.
@@
The Value Iteration algorithm is one of the main methods in dynamic programming and planning for computing the value function $V^{\pi}$ of a policy, the Policy Evaluation (PE) problem, or the optimal value function $V^{\star}$, for the Control problem. It is iteratively defined as
@@
where $V^{0}$ is the initial function. For discounted MDPs where $\gamma<1$, the Bellman operators are contractions, so by the Banach fixed-point theorem [Banach, 1922, Hunter and Nachtergaele, 2001, Hillen, 2023], the VI converges to the unique fixed points, which are $V^{\pi}$ or $V^{\star}$, with the convergence rate of $\mathcal{O}(\gamma^{k})$.
@@
If $\lambda_{1}$ is the eigenvalue such that $|\lambda_{1}|=\rho(A)$ and the other eigenvalues $\lambda_{2},\dots,\lambda_{n}$ have strictly smaller magnitude, then $(b^{k})^{\top}Ab^{k}\rightarrow\lambda_{1}$ for almost all starting points $b^{0}$ (Section 7.3.1 of Golub and Van Loan [2013]).
@@
The QR Iteration (or Orthogonal Iteration) algorithm [Golub and Van Loan, 2013, Sections 7.3 and 8.2] is a generalization of the Power Iteration that finds eigenvalues. For any $A\in\mathbb{C}^{n\times n}$, let $U^{0}\in\mathbb{C}^{n\times s}$ have orthonormal columns and perform
@@
for $k=0,1,\dotsc$. If $|\lambda_{s}|>|\lambda_{s+1}|$, the columns of $U^{k}$ converge to an orthonormal basis for the dominant $s$-dimensional invariant subspace associated with the eigenvalues $\lambda_{1},\dots,\lambda_{s}$ and diagonal entries of $(U^{s})^{\top}AU^{s}$ converge to $\lambda_{1},\dots,\lambda_{s}$ for almost all starting $U^{0}$ [Golub and Van Loan, 2013, Theorem 7.3.1].
3.  Deflated Dynamics Value Iteration
@@
3.1.  Matrix Deflation
Let $\mathcal{P}^{\pi}$ be an $n\times n$ matrix with eigenvalues $\lambda_{1},\dots,\lambda_{n}$, sorted in decreasing order of magnitude with ties broken arbitrarily. Let $u^{\top}$ denote the conjugate transpose of $u\in\mathbb{C}^{n}$. We describe three ways of deflating this matrix.
Fact 1 (Hotelling’s deflation, [Meirovitch, 1980, Section 5.6]).
@@ 3.1.  Matrix Deflation
Fact 2 (Wielandt’s deflation, [Soto and Rojo, 2006, Theorem 5]).
@@
Wielandt’s deflation, however, still requires right eigenvectors, which are sometimes numerically unstable to compute. The Schur deflation, in contrast, only requires Schur vectors, which are stable to compute  [Golub and Van Loan, 2013, Sections 7.3]. For any $\mathcal{P}^{\pi}\in\mathbb{R}^{n\times n}$, the Schur decomposition has the form $\mathcal{P}^{\pi}=URU^{\top}$, where $R$ is an upper triangular matrix with $R_{ii}=\lambda_{i}$ for $i=1,\dots,n$, and $U$ is a unitary matrix. We write $u_{i}$ to denote the $i$-th column of $U$ and call it the $i$-th Schur vector for $i=1,\dots,n$. Specifically, the QR iteration computes the top $s$ eigenvalues and Schur vectors.
Fact 3 (Schur deflation, [Saad, 2011, Proposition 4.2]).
@@
If an $E_{s}$ satisfies the conditions of Facts 1, 2, or 3, we say it is a rank-$s$ deflation matrix. Later, it will be needed that for $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$,
@@
holds. Indeed, the conditions of Facts 1, 2, or 3 do imply (1).
3.2.  Matrix Splitting: SOR
Consider the splitting of a matrix $A\in\mathbb{R}^{n\times n}$ in the form of $A=B+C+D$. Let $\alpha\neq 0$. Then, $z$ solves the linear system $Az=b$ if and only if
$(D+\alpha B)z=\alpha b-(\alpha C-(\alpha-1)D)z,$

### v2

@@ Abstract
The Value Iteration (VI) algorithm is an iterative procedure to compute the value function of a Markov decision process, and is the basis of many reinforcement learning (RL) algorithms as well. As the error convergence rate of VI as a function of iteration $k$ is $O(\gamma^{k})$, it is slow when the discount factor $\gamma$ is close to $1$. To accelerate the computation of the value function, we propose Deflated Dynamics Value Iteration (DDVI). DDVI uses matrix splitting and matrix deflation techniques to effectively remove (deflate) the top $s$ dominant eigen-structure of the transition matrix ${\mathcal{P}}^{\pi}$. We prove that this leads to a $\tilde{O}(\gamma^{k}|\lambda_{s+1}|^{k})$ convergence rate, where $\lambda_{s+1}$ is the $(s+1)$-th largest eigenvalue of the dynamics matrix. We also extend DDVI to the RL setting and present Deflated Dynamics Temporal Difference (DDTD) algorithm. We empirically show the effectiveness of the proposed algorithms.
1 Introduction
Computing the value function $V^{\pi}$ for a policy $\pi$ or the optimal value function $V^{\star}$ is an integral step of many planning and reinforcement learning (RL) algorithms. Value Iteration (VI) is a fundamental dynamic programming algorithm for computing the value functions, and its approximate and sample-based variants, such as Temporal Different Learning (Sutton, 1988), Fitted Value Iteration (Ernst et al., 2005; Munos & Szepesvári, 2008), Deep Q-Network (Mnih et al., 2015), are the workhorses of modern RL algorithms (Bertsekas & Tsitsiklis, 1996; Sutton & Barto, 2018; Szepesvári, 2010; Meyn, 2022).
@@ 1 Introduction
Recently, there has been a growing body of research that explores the application of acceleration techniques of other areas of applied math to planning and RL: Geist & Scherrer (2018); Sun et al. (2021); Ermis & Yang (2020); Park et al. (2022); Ermis et al. (2021); Shi et al. (2019) applies Anderson acceleration of fixed-point iterations, Lee & Ryu (2023) applies Anchor acceleration of minimax optimization, Vieillard et al. (2020); Goyal & Grand-Clément (2022); Grand-Clément (2021); Bowen et al. (2021); Akian et al. (2022) applies Nesterov acceleration of convex optimization, and Farahmand & Ghavamzadeh (2021); Bedaywi et al. (2024) borrow ideas from control theory/engineering, specifically PID controllers, to accelerate planning and RL algorithms.
We introduce a novel approach to accelerate VI based on modification of the eigenvalues of the transition dynamics, which are closely related to the convergence rate of VI. To see this connection, consider the policy evaluation problem where the goal is to find the value function $V^{\pi}$ for a given policy $\pi$. VI starts from an arbitrary $V^{0}$ and iteratively sets $V^{k+1}\leftarrow r^{\pi}+\gamma{\mathcal{P}}^{\pi}V^{k}$, where $r^{\pi}$ and ${\mathcal{P}}^{\pi}$ are the reward vector and the transition matrix of policy $\pi$, respectively. If $V^{0}=0$, the error vector/function at iteration $k$ is $V^{\pi}-V^{k}=\sum_{i=k}^{\infty}(\gamma{\mathcal{P}}^{\pi})^{i}r^{\pi}$. Let us take a closer look at this difference.
@@
The diagonal terms are of the form $(\gamma\lambda_{i})^{k}$, so they all converge to zero. The dominant term is $(\gamma\lambda_{1})^{k}$, which corresponds to the largest eigenvalue. As the largest eigenvalue $\lambda_{1}$ of the stochastic matrix ${\mathcal{P}}^{\pi}$ is $1$, this leads to the dominant behaviour of $O(\gamma^{k})$. This is the same rate that we also get from the cruder norm-based contraction mapping analysis. The second dominant term behaves as $O((\gamma|\lambda_{2}|)^{k})$, and so on.
If we could somehow remove from ${\mathcal{P}}^{\pi}$ the subspace corresponding to the top $s$ eigen-structure with eigenvalues $\lambda_{1},\dotsc,\lambda_{s}$, the dominant behaviour of the new procedure would be $O((\gamma|\lambda_{s+1}|)^{k})$, which can be much faster than $O(\gamma^{k})$ of the conventional VI. Although this is perhaps too good to seem feasible, this is exactly what the proposed Deflated Dynamics Value Iteration (DDVI) algorithm achieves. Even if $\mathcal{P}^{\pi}$ is not a diagonalizable matrix, DDVI works. DDVI is based on two main ideas.
The first idea is the deflation technique, well-studied in linear algebra (see Section 4.2 of Saad 2011), that allows us to remove large eigenvalues of ${\mathcal{P}}^{\pi}$. This is done by subtracting a matrix $E$ from ${\mathcal{P}}^{\pi}$. This gives us a “deflated dynamics” ${\mathcal{P}}^{\pi}-E$ that does not have eigenvalues $\lambda_{1},\dotsc,\lambda_{s}$. The second idea is based on matrix splitting (see Section 11.2 of Golub & Van Loan 2013), which allows us to use the modified dynamics ${\mathcal{P}}^{\pi}-E$ to define a VI-like iterative procedure and still converge to the same solution $V^{\pi}$ to which the conventional VI converges.
Deflation has been applied to various algorithms such as conjugate gradient algorithm (Saad et al., 2000), principal component analysis (Mackey, 2008), generalized minimal residual method (Morgan, 1995), and nonlinear fixed point iteration (Shroff & Keller, 1993) for improvement of convergence. While some prior work in accelerated planning can be considered as special cases of deflation (Bertsekas, 1995; White, 1963), to the best of our knowledge, this is the first application of the deflation technique that eliminates multiple eigenvalues in the context of RL. On the other hand, multiple planning algorithms have been introduced based on the general idea of matrix splitting (Hastings, 1968; Kushner & Kleinman, 1968; 1971; Reetz, 1973; Porteus, 1975; Rakhsha et al., 2022), though with a different splitting than this work.
@@
2 Background
We first briefly review basic definitions and concepts of Markov Decision Processes (MDP) and Reinforcement Learning (RL) (Bertsekas & Tsitsiklis, 1996; Sutton & Barto, 2018; Szepesvári, 2010; Meyn, 2022). We then describe the Power Iteration and QR Iterations, which can be used to compute the eigenvalues of a matrix.
@@ 2 Background
The discounted MDP is defined by the tuple $(\mathcal{X},\mathcal{A},\mathcal{P},\mathcal{R},\gamma)$, where $\mathcal{X}$ is the state space, $\mathcal{A}$ is the action space, $\mathcal{P}\colon\mathcal{X}\times\mathcal{A}\rightarrow\mathcal{M}(\mathcal{X})$ is the transition probability kernel, $\mathcal{R}\colon\mathcal{X}\times\mathcal{A}\rightarrow\mathcal{M}()$ is the reward kernel, and $\gamma\in[0,1)$ is the discount factor. In this work, we assume that the MDP has a finite number of states. We use $r\colon\mathcal{X}\times\mathcal{A}\rightarrow$ to denote the expected reward at a given state-action pair. Denote $\pi\colon\mathcal{X}\rightarrow\mathcal{M}(\mathcal{A})$ for a policy. We define the reward function for a policy $\pi$ as $r^{\pi}(x)=\mathbf{E}_{a\sim\pi(\cdot\mid x)}\left[r(x,a)\right]$. The transition kernel of following policy $\pi$ is denoted by ${\mathcal{P}}^{\pi}\colon\mathcal{X}\rightarrow\mathcal{M}(\mathcal{X})$ and is ${\mathcal{P}}^{\pi}(x^{\prime}\mid x)=\sum_{a\in\mathcal{A}}\pi(a\mid x)\mathcal{P}(x^{\prime}\mid x,a)$ (or an integral over actions, if the action space is continuous).
The value function for a policy $\pi$ is $V^{\pi}(x)=\mathbf{E}_{\pi}[\sum^{\infty}_{t=0}\gamma^{t}r(x_{t},a_{t})\mid x_{0}=x]$ where $\mathbf{E}_{\pi}$ denotes the expected value over all trajectories $(x_{0},a_{0},x_{1},a_{1},\dots)$ induced by $\mathcal{P}$ and $\pi$. We say that $V^{\star}$ is optimal value functions if $V^{\star}=\sup_{\pi}V^{\pi}$. We say $\pi^{\star}$ is an optimal policy if $\pi^{\star}\in\operatorname*{Arg\,max}_{\pi}{V^{\pi}}$.
The Bellman operator $T^{\pi}$ for policy $\pi$ is defined as the mapping that takes $V\in\mathcal{F}(\mathcal{X})$ and returns a new function such that its value at state $x$ is $(T^{\pi}V)(x)=\mathbf{E}_{a\sim\pi(\cdot\mid x),x^{\prime}\sim\mathcal{P}(\cdot\mid x,a)}\left[r(x,a)+\gamma V(x^{\prime})\right]$. The Bellman optimality operator ${T^{\star}}$ is defined as $({T^{\star}}V)(x)=\sup_{a\in\mathcal{A}}\left\{r(x,a)+\gamma\mathbf{E}_{x^{\prime}\sim\mathcal{P}(\cdot\,|\,x,a)}\left[V(x^{\prime})\right]\right\}$. The value functions $V^{\pi}$ and $V^{\star}$ are the fixed points of the Bellman operators, that is, $V^{\pi}=T^{\pi}V^{\pi}$ and $V^{\star}=T^{\pi}V^{\star}$.
@@
The Value Iteration algorithm is one of the main methods in dynamic programming and planning for computing the value function $V^{\pi}$ of a policy (the Policy Evaluation (PE) problem), or the optimal value function $V^{\star}$ (the Control problem). It is iteratively defined as
@@
where $V^{0}$ is the initial function. For discounted MDPs where $\gamma<1$, the Bellman operators are contractions, so by the Banach fixed-point theorem (Banach, 1922; Hunter & Nachtergaele, 2001; Hillen, 2023), the VI converges to the unique fixed points, which are $V^{\pi}$ or $V^{\star}$, with the convergence rate of $O(\gamma^{k})$.
@@
If $\lambda_{1}$ is the eigenvalue such that $|\lambda_{1}|=\rho(A)$ and the other eigenvalues $\lambda_{2},\dots,\lambda_{n}$ have strictly smaller magnitude, then $(b^{k})^{\top}Ab^{k}\rightarrow\lambda_{1}$ for almost all starting points $b^{0}$ (Section 7.3.1 of Golub & Van Loan 2013).
@@
The QR Iteration (or Orthogonal Iteration) algorithm (Golub & Van Loan, 2013, Sections 7.3 and 8.2) is a generalization of the Power Iteration for finding multiple eigenvalues. For any $A\in\mathbb{C}^{n\times n}$, let $U^{0}\in\mathbb{C}^{n\times s}$ have orthonormal columns and perform
@@
for $k=0,1,\dotsc$. If $|\lambda_{s}|>|\lambda_{s+1}|$, the columns of $U^{k}$ converge to an orthonormal basis for the dominant $s$-dimensional invariant subspace associated with the eigenvalues $\lambda_{1},\dots,\lambda_{s}$ and diagonal entries of $(U^{s})^{\top}AU^{s}$ converge to $\lambda_{1},\dots,\lambda_{s}$ for almost all starting $U^{0}$ (Golub & Van Loan, 2013, Theorem 7.3.1). Each QR iteration needs $O(n^{2}s)$ flops for its matrix multiplication, which is the dominant part of the computation, and $O(ns^{2})$ flops for the QR factorization.
3 Deflated Dynamics Value Iteration
@@
3.1 Matrix Deflation
Recall from the introduction that we would like to remove from ${\mathcal{P}}^{\pi}$ the subspace corresponding to the top $s$ eigen-structure with eigenvalues $\lambda_{1},\dotsc,\lambda_{s}$. We use the so-called deflation technique to displace, and in fact remove, some of the eigenvalues of the matrix ${\mathcal{P}}^{\pi}$ without changing the rest. This is done by subtracting a matrix $E$ of a certain form from ${\mathcal{P}}^{\pi}$ (see Section 4.2 of Saad 2011 for a discussion of deflation technique).
Let $\mathcal{P}^{\pi}$ be an $n\times n$ matrix with eigenvalues $\lambda_{1},\dots,\lambda_{n}$, sorted in decreasing order of magnitude with ties broken arbitrarily. Let $u^{\top}$ denote the conjugate transpose of $u\in\mathbb{C}^{n}$. We describe three ways of deflating this matrix and some properties of each approach as the following Facts.
Fact 1 (Hotelling’s deflation, Meirovitch 1980, Section 5.6).
@@ 3.1 Matrix Deflation
Fact 2 (Wielandt’s deflation, Soto & Rojo 2006, Theorem 5).
@@
Wielandt’s deflation, however, still requires right eigenvectors, which are sometimes numerically unstable to compute. The Schur deflation, in contrast, only requires Schur vectors, which are stable to compute  (Golub & Van Loan, 2013, Sections 7.3). For any $\mathcal{P}^{\pi}\in\mathbb{R}^{n\times n}$, the Schur decomposition has the form $\mathcal{P}^{\pi}=URU^{\top}$, where $R$ is an upper triangular matrix with $R_{ii}=\lambda_{i}$ for $i=1,\dots,n$, and $U$ is a unitary matrix. We write $u_{i}$ to denote the $i$-th column of $U$ and call it the $i$-th Schur vector for $i=1,\dots,n$. Specifically, the QR iteration computes the top $s$ eigenvalues and Schur vectors.
Fact 3 (Schur deflation, Saad 2011, Proposition 4.2).
@@
If an $E_{s}$ satisfies the conditions of Facts 1, 2, or 3, we say it is a rank-$s$ deflation matrix. Later, it will be needed that for $E_{s}=\sum_{i=1}^{s}\lambda_{i}u_{i}v_{i}^{\top}$, the identity
@@
should hold. Indeed, the conditions of Facts 1, 2, or 3 do imply equation 1.
3.2 Matrix Splitting: Successive over-relaxation
We now describe the second key ingredient of DDVI: matrix splitting. Recall that the policy evaluation problem is the problem of finding the value function $V^{\pi}$ such that $T^{\pi}V^{\pi}=V^{\pi}$. This is in fact a system of linear equations:
$(I-\gamma\mathcal{P}^{\pi})V^{\pi}=r^{\pi}.$ (2)
DDVI’s approach to solve this equation is based on the Successive over-relaxation (SOR) algorithm, an iterative approach to solve linear systems.
To introduce SOR in its general form, consider a matrix $A\in\mathbb{R}^{n\times n}$ and the linear system $Az=b$. SOR starts by splitting $A$ in the form of $A=B+C+D$. Let $\alpha\neq 0$. Then, $z$ solves the linear system $Az=b$ if and only if
$(D+\alpha B)z=\alpha b-(\alpha C+(\alpha-1)D)z,$